Matrices multiplication and Partitioned multiplication:

matrix X=

2 1 5
4 2 3

Matrix Y=

1 2 4
2 3 1

1. Find the XY^(T) T means transpose

2.Compute the outer product expansion of XY^(T) .

3. did you get the same answer from 1 and 2?
Answer #1

(1)
XYT is given by:

4 2 3)×123

c11 = (2X1) + (1 X2)+(5X4) = 24

c12=(2X2)+(1X3)+(5X1)=12

c21=(4X1)+(2X2)+(3X4)=20

c22=(4X2)+(2X3)+(3X1)=17

So,

Answer is:

C=XY^{T}=egin{pmatrix} 24 &12 20 & 17 end{pmatrix}

(2)

The outer product expansion of XYT is

061 243612 486 12 1953 2 530 10 6 20 2448 244 8.8

(3)

We did not get the same answer from 1 and 2.